Demand Management-situation assessment

Consider the following information.

You have a MTS product that has sales of about 1000 units per month. You manufacture this product in 500 unit batches (so, roughly, a 500 unit batch every other week). There is no significant seasonality and the current standard deviation is 30 units per month.

Customer orders: The average order size is 15 units; the largest, single order within any month is usually no more than 30. Orders are received daily and more or less evenly distributed throughout the month. Orders are normally shipped within 24 hours of receipt (consistent with operating in a MTS environment).

As of the start of business this morning (a Monday), you have 330 units in inventory, and the next batch of 500 units is due in next Monday. You operate on a five-day-per-week (M-F) schedule.

A single order for 150 units has just been received, with no special shipping instructions or ship dates (implying that normal order turn-around time is expected?). Would you want your order entry system to let this order into your system for fulfillment just like any other order? Yes or no and briefly explain your reasoning. What other actions should you then take?
Answer #1

Monthly demand 1000 units

Working days 20 days I.e(M-F) or 22 days

Average demand per day 1000/20=50 units

Average demand per day 1000/22=45.45 units or 45 units

Opening stock on monday 330 units

If we enter 150 units in the system balance units is 330-150= 180 units

Production is on next monday working days in between is 5 days i.e M-F

And no other demand on monday except 150 units

If we take less average i.e 45 days then After tuesday balance units 180-45= 135 units

after Wednesday 135-45=90

After thursday 90-45=45

After friday 45-45=0

Conclusion:- if we consider the above situation 150unitsr is acceptable with no order on moday with 45 units average daily demand.

if demand is more than 45 units then we can not enter the 150 units order into the system

if demand on monday is also for 45 units or 50 units + 150 units then we can not enter into the system.